OpenStudy (anonymous):

Simplify. 10x - 3y + 2x - 4y

OpenStudy (anonymous):

x and y are our variables, so we can "pair" them up, so to speak. 10x +2x= 12x -3y -4y= -7y 12x -7y

OpenStudy (anonymous):

Evaluate and simplify if possible: 3(x - 2y) - 4(2x + 3y) for x = -1 and y = 3

OpenStudy (anonymous):

ok any time soon

OpenStudy (anonymous):

3( -1 -2(3)) - 4)(2(-1)+3(3)) 3(-1 -6) -4 ( -2 +9) 3( -7) - 4(7) -21 - 28 -49

OpenStudy (anonymous):

\[10x-3y+4y-2x\]